OpenStudy (anonymous):

Solve. Express your answer in simplest form. 2 1/4÷ 3 3/4

OpenStudy (saifoo.khan):

Do you know how to convert mixed fractions into improper?

OpenStudy (anonymous):

it's like this 2 1/4÷ 3 3/4 -> 6/4 x 4/15

OpenStudy (saifoo.khan):

Get this?\[2\frac14 = \frac{(2\times 4)+1}{4}= \frac94\]

OpenStudy (anonymous):

so the answer is 3/5

OpenStudy (anonymous):

oh so 9/4

OpenStudy (anonymous):

ok thanks

OpenStudy (saifoo.khan):

Now convert 3 3/4 in this form.